• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ipom

    Log entry time 08:10:48 on March 26, 2011

    Entry number 353274

    keyword=Owl shift summary

    00:00 Shift starts, continue taking 4he SRC production data
    00:02 Stopped run 3052 to change CER HV (yez).
    01:25 Test run on 8uA, ps1=160,ps2=1,ps3=1,ps4=1 we get T1=70Hz, T3=220Hz  
    DT=10%
    02:40 Test run on 2uA, ps1=40,ps2=1,ps3=1,ps4=1 we get T1=80Hz, T3=71Hz  
    DT=8%
    03:40 back to 4uA, ps1=80,ps2=1,ps3=1,ps4=1 we get T1=70Hz, T3=130Hz  
    DT=9%
    
    Total collected charge this shift 3052-3061 :  0.12288 c
    
    
    3052 4.5uA, 0.79M events, charge: 0.01650c
    3053 4.5uA, 1M events, charge: 0.02147 c
    3054 JUNK
    3055 8uA,   1M events, charge: 0.02917 c
    3056 2uA,   0.5M events. charge: 0.007198 c
    3057 JUNK
    3058 2uA,   1M events. charge: 0.02027 c
    3059 2uA,   1M events. charge: 0.02167 c
    3060 JUNK - coda crushed
    3061 2uA,   300K events, 0.005686 c